A phase change has ASsys = -17.250x10^1) J/K mol and AHsys = -(3.55x10^1) kJ/mol. At what temperature (in K) will this phasePlease help is due in 30 mins, will rate, Thank you!
Answer #1

We know the following equation:

\Delta G=\Delta H-T\Delta S

Now, at equilibrium change in free energy becomes zero (\Delta G=0)

Therefore, \Delta H=T\Delta S

\frac{\Delta H}{\Delta S}=T

\frac{-3.55*10^4(J/mol)}{-7.25*10^1(J/molK)}=T

489.65K=T